# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v2

@@ 4.1 Deformation to a product metric and stability
    For $\varepsilon>0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{\varepsilon})=\{0\}$. For $\varepsilon=0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{0})=\mathbb{R}d\phi_{\pm}$.
@@
  4. 4.
    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon\in[0,1]$.
@@
(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$. (2) For $\varepsilon>0$, Lemma 2.4 shows $d\phi_{\pm}$ are not parallel; the proof of Lemma 4.5 will show no other parallel 1-forms exist. For $\varepsilon=0$, $d\phi_{\pm}$ are clearly parallel. (3) is Lemma 4.6 (metric invariance).
(4) For $\varepsilon=0$, the harmonic representative of $\omega]$ is
$\omega_{h}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}.$
The subspace $\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$ consists of those harmonic 3‑forms that can be written as
$\alpha\wedge\beta,$
where $\alpha$ is a parallel 2‑form on $T^{2}$ (hence a constant multiple of $\text{vol}_{\text{BZ}}$) and $\beta$ is a parallel 1‑form on the fibres (hence a constant linear combination of $d\phi_{+}$ and $d\phi_{-}$).
If $\omega_{h}$ belonged to this subspace, we would have
$F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\alpha\wedge\beta=(\lambda\text{vol}_{\text{BZ}})\wedge(c_{+}d\phi_{+}+c_{-}d\phi_{-}),$
which would imply $F^{(+)}=\lambda c_{+}\text{vol}_{\text{BZ}}$ and $F^{(-)}=\lambda c_{-}\text{vol}_{\text{BZ}}$ as differential forms, not just in cohomology. In a generic SOC texture, $F^{(+)}$ and $F^{(-)}$ are not parallel (they are not constant multiples of $\text{vol}_{\text{BZ}}$ pointwise), hence $\omega_{h}\notin\mathcal{P}_{2}\otimes\mathcal{P}_{1}$. Therefore $\mathcal{K}_{0}=0$. ∎
@@
By Lemma 4.2, the topological data $r^{\sharp}=r-\dim\mathcal{K}_{\varepsilon}$ is constant (=r=r) for all $\varepsilon$. The inequality holds at $\varepsilon=0$ by assumption. By Lemma 4.3, the set $\{\varepsilon:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}\}$ is closed (by lower semi-continuity). Since it contains $\varepsilon=0$ and $[0,1]$ is connected, it must be all of $[0,1]$. ∎
4.2 Parallel forms in Kaluza–Klein metrics
Lemma 4.5 (Parallel forms in Kaluza–Klein metrics).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
  1. 1.
    If all connection forms $A^{a}$ are closed (i.e., $dA^{a}=0$), then every $h$-parallel form on the fibre is also $g$-parallel.
  2. 2.
    If at least one curvature $F^{a}:=dA^{a}$ is not identically zero, then the space of $g$-parallel 1-forms tangent to the fibres is trivial:
    $\mathcal{P}_{1}(M_{2},g)=\{0\}.$
Proof.
Write the metric locally as $g=g_{ij}(x)dx^{i}dx^{j}+\sum_{a=1}^{m}(\Theta^{a})^{2}$, where $\Theta^{a}=\theta^{a}+A^{a}_{i}(x)dx^{i}$ and $\{\theta^{a}\}$ is the orthonormal coframe on $M_{2}$ induced by the flat metric $h$. In our SOC BEC case, $m=2$, $\theta^{1}=d\phi_{+}$, $\theta^{2}=d\phi_{-}$.
Let $\xi=\sum_{a=1}^{m}f_{a}(x,y)\Theta^{a}$ be a 1‑form tangent to the fibres. The Levi‑Civita connection of $g$ satisfies the following relation for a horizontal vector field $X=X^{i}\partial_{i}$ and the vertical coordinate field $V_{a}$ dual to $\theta^{a}$:
$\nabla^{LC}_{X}V_{a}=\tfrac{1}{2}F^{a}(X)^{\#},$
where $F^{a}(X)^{\#}$ is the vertical vector field obtained by raising the index of the 1‑form $F^{a}(X,\cdot)$ with the fibre metric $h$. Formula (i) follows from the Koszul formula and uses only that the fibres are flat.
Now suppose $\xi$ is parallel, $\nabla^{LC}\xi=0$. Evaluating on the pair $(X,V_{a})$ gives
$0=(\nabla^{LC}_{X}\xi)(V_{a})=X\bigl(\xi(V_{a})\bigr)-\xi(\nabla^{LC}_{X}V_{a})=X(f_{a})-\tfrac{1}{2}\sum_{b=1}^{m}f_{b}\,h\!\bigl(F^{a}(X)^{\#},V_{b}\bigr).$
Because the functions $f_{a}$ are constant along the fibres (they depend only on the base coordinates $x$ due to the left‑invariance of the metric $h$), equation (ii) reduces to
$\tfrac{1}{2}\sum_{b=1}^{m}f_{b}(x)\,F^{b}(X,V_{a})=0\qquad\text{for every horizontal }X\text{ and every }a.$
If all $F^{b}=0$, condition (iii) is vacuous and any choice of constant $f_{b}$ yields a parallel form, proving statement (i). If, however, some $F^{b_{0}}$ is not identically zero, we can choose a horizontal vector field $X$ and an index $a$ such that $F^{b_{0}}(X,V_{a})\neq 0$ at some point; then (iii) forces $f_{b_{0}}=0$. Repeating the argument for every independent curvature forces all coefficients $f_{b}$ to vanish, hence $\xi=0$. This proves statement (ii) for 1‑forms.
The argument for higher‑degree forms is analogous, using that a parallel $k$-form on the fibre must have constant coefficients with respect to the left‑invariant coframe, and the same curvature coupling appears in the covariant derivative. ∎
Lemma 4.6 (Harmonic representative and Künneth component).
Let $g$ be a Kaluza–Klein metric as in Definition 4.1 on $M=M_{1}\times M_{2}$. Let $\omega]\in H^{k}(M;\mathbb{R})$ be a cohomology class and $\omega_{h}$ its unique $g$-harmonic representative. Then $\omega_{h}$ respects the Künneth decomposition in the sense that its projection onto the $(p,q)$-bidegree component corresponds to the harmonic representative of the mixed class $\omega]_{p,q}\in H^{p}(M_{1})\otimes H^{q}(M_{2})$. In particular, $\omega_{h}$ admits a unique decomposition
$\omega_{h}=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i}+\gamma,$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$, the forms $\alpha_{i},\beta_{i}$ are the lifted harmonic representatives of the factor classes, and $\gamma$ contains terms of different bidegrees.
Proof.
By Hodge theory on compact Riemannian manifolds, every de Rham cohomology class $\omega]$ possesses a unique harmonic representative $\omega_{h}$ with respect to the metric $g$. The Künneth theorem provides a canonical isomorphism $H^{k}(M)\cong\bigoplus_{p+q=k}H^{p}(M_{1})\otimes H^{q}(M_{2})$.
While the Hodge Laplacian $\Delta_{g}$ of a Kaluza-Klein metric does not generally commute with the factor-wise Laplacians $\Delta_{g_{1}}$ and $\Delta_{h}$ due to the presence of the connection forms $A^{a}$, the existence of the harmonic representative for each mixed component is guaranteed by the ellipticity of $\Delta_{g}$. The tensor rank $r$ is a purely algebraic property of the class $\omega]$ in the tensor product of the cohomology groups $H^{*}(M_{i})$ and is therefore independent of the specific metric $g$.
The decomposition follows by projecting the global harmonic form $\omega_{h}$ onto the subspace of forms of pure bidegree, which remains well-defined under the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the Kaluza-Klein structure. ∎
@@ 4.5 The PT bound for Kaluza–Klein metrics with flat fibres
Theorem 4.10 (PT bound for Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold equipped with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $\{\theta^{a}\}$ defines a flat left‑invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$. Let $\nabla^{C}=\nabla^{LC}(g)+\frac{1}{2}T$ be a metric connection on $(M,g)$ with totally skew‑symmetric torsion $T\in\Omega^{3}(M)$. Denote by $\omega_{h}$ the unique $g$-harmonic representative of $T]\in H^{3}(M;\mathbb{R})$ and assume
  1. 1.
    $\nabla^{LC}(g)T\neq 0$;
  2. 2.
    $\omega_{h}]$ has a non‑trivial mixed component in the Künneth decomposition;
  3. 3.
    the mixed tensor rank $r:=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is finite.
Define the obstruction space
$\mathcal{K}\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(\mathcal{P}_{2}(M_{1},g)\otimes\mathcal{P}_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(\mathcal{P}_{1}(M_{1},g)\otimes\mathcal{P}_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim\mathcal{K}$.
Then for every point $p\in M$,
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.0)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $\mathcal{P}_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
Proof.
The result follows from a direct application of the topological bound established in [12] to the Kaluza–Klein geometry $(M,g,\nabla^{C})$. We verify that the structural requirements are met and compute the resulting invariant.
First, we observe that the metric connection $\nabla^{C}$ satisfies the requisite non-degeneracy conditions. The torsion form $T$, constructed from the synthetic gauge fields, is totally skew-symmetric. Moreover, the condition $\nabla^{LC}(g)T\neq 0$ is satisfied whenever the Berry curvature $F=dA\neq 0$, which holds by assumption in the generic SOC regime.
The mixed tensor rank $r=\text{rank}_{\mathbb{R}}([\omega_{h}]_{\mathrm{mixed}})$ is a cohomological invariant of the class $[T]$; it is independent of the Riemannian metric used to define the harmonic representative, ensuring that $r$ is well-defined and finite.
Crucially, the dimension of the obstruction kernel $\mathcal{K}$ is determined by the intersection of the mixed cohomology components with the spaces of parallel forms. By Lemma 4.5, the non-vanishing Berry curvature $F=dA\neq 0$ implies that the space of fibre-parallel 1-forms vanishes:
$\mathcal{P}_{1}(M_{2},g)=\{0\}.$
Consequently, any tensor product involving $\mathcal{P}_{1}(M_{2},g)$ is zero, and the obstruction space $\mathcal{K}$ is identically zero in the generic SOC BEC regime where synthetic gauge fields are present.
Thus, we have $\dim\mathcal{K}=0$, and the general theorem of [12] then states that the dimension of the off-diagonal holonomy is bounded by $r-\dim\mathcal{K}$, yielding
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r.$
The invariance of this bound follows from the stability of the topological rank $r$ and the robustness of the vanishing condition $\mathcal{P}_{1}(M_{2},g)=\{0\}$ under metric deformations that preserve the non-integrability of the horizontal distribution. ∎

### v3

@@ 4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
The general topological lower bound of Pigazzini and Toda [12] is stated for genuine Riemannian product manifolds. Our physical metric (2.1), however, is a Kaluza–Klein metric that does not globally factor as a product. We now explain why, despite this apparent mismatch, the bound applies to our specific geometry. The argument relies on three observations that hold for the particular Kaluza–Klein structure arising from a two–component SOC BEC.
Cohomology class and its harmonic representative
The torsion 3–form $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$ need not be closed, but its de Rham class $\omega]=[T^{b}]\in H^{3}(M;\mathbb{R})$ is well defined. (Recall that $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$; note that the pullback $\pi^{*}\bigl(F^{(\pm)}\wedge A^{(\pm)}\bigr)$ vanishes because on the 2-dimensional base $T^{2}_{\mathrm{BZ}}$ we have $F^{(\pm)}\wedge A^{(\pm)}=0$. Therefore, $T^{b}$ is cohomologous to $F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}$.) Under the Künneth decomposition
$H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
the class $\omega]$ can be written as
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry–curvature classes are necessarily proportional; hence there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$. Consequently,
which is a single simple tensor. The mixed tensor rank is therefore $r=1$, a purely topological invariant.
Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of $\omega]$. Under Assumption 2.1, Lemma 3.1 gives an explicit form:
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
where $a_{\pm}$ are constants with $a_{-}=\lambda a_{+}$ and $\lambda=c^{(-)}/c^{(+)}$. This factorised form is the exact analogue, for our Kaluza–Klein metric, of the factorised harmonic decomposition used in the product-metric case.
Spaces of parallel forms
The bound involves the obstruction kernel
$\mathcal{K}=\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr).$
For the flat base we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}$. By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}=0$.
Pointwise construction of the obstruction maps
Even though the metric is not a product, at each point $p$ we have an orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. The non-integrability of $\mathcal{H}$ (measured by the Berry curvatures $F^{(\pm)}$) does not affect pointwise evaluations. Consider the (2,1)–component. Using the explicit harmonic representative from Lemma 3.1, we set
$\mathcal{V}_{2,1}:=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\}\subset H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}).$
At a point $p=(k,\phi_{+},\phi_{-})$ choose any vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and a vertical vector $Z\in\mathcal{V}_{p}$ such that $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$. Define
$\Psi_{p}\bigl(\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\bigr):=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
where $(\cdot)_{\mathrm{off}}$ denotes the part mapping $\mathcal{H}$ to $\mathcal{V}$ or vice versa. The map is well-defined because any other choice $(X^{\prime},Y^{\prime},Z^{\prime})$ satisfying the same conditions gives the same operator up to a constant factor, due to the linearity of the curvature in its arguments.
A direct computation (provided in Appendix B) shows that for $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, we have:
$\Psi_{p}(\xi)=\frac{1}{2}\left(a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))\right)J,$
where $J:T_{p}M\to T_{p}M$ is a fixed off-diagonal operator (independent of $a_{\pm}$) that exchanges horizontal and vertical directions. The calculation uses:
  1. 1.
    $\nabla^{LC}(\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)})=\operatorname{vol}_{\mathrm{BZ}}\wedge\pi^{*}F^{(\pm)}$ (since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base);
  2. 2.
    the definition of $\Psi_{p}$ in terms of $(\nabla^{LC}T)_{\mathrm{off}}$;
  3. 3.
    the choice of vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and $Z\in\mathcal{V}_{p}$ with $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$.
Therefore, $\Psi_{p}(\xi)=0$ if and only if $a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))=0$. Under Assumption 2.1, $F^{(-)}=\lambda F^{(+)}$ with $\lambda=c^{(-)}/c^{(+)}$. Moreover, since $\xi\in\mathcal{V}_{2,1}$, the coefficients satisfy $a_{-}=\lambda a_{+}$ (because $\mathcal{V}_{2,1}$ is spanned by $\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})$). Substituting gives
$a_{+}F^{(+)}+a_{-}F^{(-)}=a_{+}F^{(+)}+\lambda a_{+}(\lambda F^{(+)})=a_{+}(1+\lambda^{2})F^{(+)}.$
Since $F^{(+)}\neq 0$, this vanishes precisely when $a_{+}=0$, which also implies $a_{-}=0$. Hence $\ker\Psi_{p}=\{0\}$ and $\operatorname{im}\Psi_{p}$ is one-dimensional. The same argument applied to the (1,2)–component gives a map $\widetilde{\Psi}_{p}$ with trivial kernel.
Because the images of $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ consist of off–diagonal curvature operators, the Ambrose–Singer theorem forces
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\dim\operatorname{im}\Psi_{p}+\dim\operatorname{im}\widetilde{\Psi}_{p}=1+0=1.$
Thus the topological lower bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The deformation family introduced below will serve only to connect this result to the product–metric case and to make the semicontinuity argument transparent.
Note 1: The reasoning above does not require the metric to be a global Riemannian product. It uses only the pointwise orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$, the explicit description of the harmonic representative (Lemma 3.1, which relies on the flatness of the base and the fibres under Assumption 2.1), and the vanishing of $\mathcal{P}_{1}$ on the fibres (Lemma 2.5). All these ingredients are present in the SOC BEC geometry.
Deformation to a product metric and stability
To provide a transparent link with the original product–metric setting, and to control the behaviour of the holonomy dimension via lower–semicontinuity, we now introduce a smooth deformation that connects the physical Kaluza–Klein metric to a genuine product metric.
Note 2: [On the wedge decomposition in Kaluza–Klein geometry] In the product case, the harmonic representative decomposes as a sum of wedge products of harmonic forms on the factors. In our Kaluza–Klein geometry, the analogous decomposition (Lemma 3.1) involves the horizontal forms $\Theta^{(\pm)}$, which are harmonic with respect to the full metric $g_{M}$ but are not pullbacks from the fibres alone. However, they satisfy:
$\Delta_{g_{M}}\Theta^{(\pm)}=0\quad\text{and}\quad[\Theta^{(\pm)}]=[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}),$
which is sufficient for the cohomological arguments. The key point is that $\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)}$ represents the mixed class $F^{(\pm)}]\otimes[d\phi_{\pm}]$ in the Künneth decomposition, regardless of the non-product nature of the metric.
@@ 4.1 Deformation to a product metric and stability
    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies:
    • (a)
      $\dim\mathcal{K}_{0}=1$.
    • (b)
      $\dim\mathcal{K}_{\varepsilon}=0$ for every $\varepsilon>0$.
    Consequently, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$.
@@
(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$.
(2) The mixed tensor rank is a topological invariant (Lemma 4.7) and therefore constant under metric deformations.
(3a) For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, the Brillouin torus is flat. On a flat Riemannian manifold every harmonic form is parallel; thus $\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$. The harmonic representative of the torsion class $\omega]$ with respect to $g_{0}$ is
$\omega_{h}^{(0)}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-},$
where $\alpha_{\pm}$ are the harmonic representatives of the Berry‑curvature classes $F^{(\pm)}]$. Since $\alpha_{\pm}$ are parallel and $d\phi_{\pm}$ are parallel, $\omega_{h}^{(0)}$ belongs to $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$. Hence $\mathcal{K}_{0}=\mathcal{V}_{2,1}$ and $\dim\mathcal{K}_{0}=1$.
(3b) For $\varepsilon>0$ the metric $g_{\varepsilon}$ is a Kaluza–Klein metric satisfying Assumption 2.1 (with curvatures scaled by $\varepsilon$). By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$.
Thus $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$. ∎
@@
By Lemma 4.2, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ is constant on $(0,1]$ and equals $1$. For every $\varepsilon>0$, Theorem 4.8 (the extension of the PT theorem to Kaluza–Klein metrics) therefore yields the non‑trivial lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$.
The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower‑semicontinuous by Lemma 4.3; hence the set
$S=\{\varepsilon\in[0,1]:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\}$
is closed. Since $S$ contains the interval $(0,1]$, its closure contains the endpoints $\varepsilon=0$ and $\varepsilon=1$. Because $S$ is closed, we must have $0\in S$ and $1\in S$. In particular, $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$, which is the desired bound for the physical metric. ∎
@@ 4.4 The PT bound for Kaluza–Klein metrics with flat fibres
Theorem 4.8 (PT bound for the SOC BEC geometry).
Let $M=T_{\mathrm{BZ}}^{2}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be equipped with the Kaluza-Klein metric $g_{M}$ as in (2.1), and let $\nabla^{C}=\nabla^{LC}+\frac{1}{2}T$ be the metric connection with torsion $T=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$. Assume that $\nabla^{LC}T\neq 0$ and that the Berry curvature classes satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$. Then the off-diagonal holonomy of $\nabla^{C}$ satisfies
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq 1\quad\text{for every }p\in M.$
Proof for the SOC BEC geometry.
We prove the bound via deformation and lower semi-continuity, without requiring global integrability of the horizontal distribution.
Consider the family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1. For each $\varepsilon>0$, the orthogonal splitting $T_{p}M=\mathcal{H}^{(\varepsilon)}_{p}\oplus\mathcal{V}_{p}$ (horizontal/vertical) is pointwise well-defined. Although for $\varepsilon>0$ this splitting is not integrable unless $F^{(\pm)}=0$, the original Pigazzini–Toda theorem [12] uses only the pointwise orthogonal decomposition of the tangent space, not its integrability. Indeed, the proof of Theorem 5.2 in [12] relies solely on algebraic computations with curvature operators and on the properties of harmonic forms, which remain valid for any orthogonal splitting.
For $\varepsilon>0$, Lemma 2.5 implies that the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$. The mixed tensor rank $r=1$ is topological (Lemma 4.7), so $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}=1$.
Applying Theorem 5.2 of [12] to the splitting $\mathcal{H}^{(\varepsilon)}\oplus\mathcal{V}$ yields
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}_{\varepsilon}=1\quad\forall\varepsilon>0,$
where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$. This application is justified because Theorem 5.2 of [12] requires only a pointwise orthogonal decomposition $T_{p}M=V_{1}\oplus V_{2}$ and the existence of the harmonic representative $\omega_{h}$ with the stated properties. The non-integrability of $\mathcal{H}^{(\varepsilon)}$ (when $\varepsilon>0$) does not affect the proof, which is based on pointwise algebraic computations with curvature operators.
Define $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ for $\varepsilon\in[0,1]$. Then $F$ is integer-valued (a dimension) and, by Lemma 4.3, lower semi-continuous on $[0,1]$.
We now prove $F(1)\geq 1$. For an integer-valued lower semi-continuous function, the set $U:=\{\varepsilon\in[0,1]\mid F(\varepsilon)\geq 1\}$ is open in $[0,1]$: if $\varepsilon_{0}\in U$, lower semi-continuity gives $\delta>0$ such that $F(\varepsilon)>F(\varepsilon_{0})-\frac{1}{2}$ for all $\varepsilon$ in $(\varepsilon_{0}-\delta,\varepsilon_{0}+\delta)\cap[0,1]$. Since $F(\varepsilon_{0})\geq 1$ and $F$ is integer, this implies $F(\varepsilon)\geq 1$ on that neighbourhood, hence the neighbourhood is contained in $U$.
We have shown $F(\varepsilon)\geq 1$ for all $\varepsilon>0$; thus $(0,1]\subseteq U$. Because $U$ is open and contains $(0,1]$, any open neighbourhood of $1$ in $[0,1]$ intersects $(0,1)\subset U$; openness of $U$ then forces $1\in U$. Therefore $F(1)\geq 1$, i.e., $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$.
Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, the bound holds for the physical metric. ∎